40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Jul 29 22:37:54 2022, Koji, Update, IOO, WFS investigation PXL_20220730_040900843.jpgPXL_20220730_041216848.jpg
    Reply  Tue Aug 2 01:10:26 2022, Koji, Update, IOO, WFS investigation 220801_IMC_WFS_DEMOD.pdf220801_IMC_WFS_DEMOD2.pdfWFS1.pngWFS2.png
       Reply  Thu Aug 4 21:58:18 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-10-57.pngScreenshot_2022-08-04_22-11-16.pngScreenshot_2022-08-04_22-11-53.pngScreenshot_2022-08-04_22-12-39.png
          Reply  Thu Aug 4 22:14:20 2022, Koji, Update, IOO, WFS investigation IMC_SUS_channels_TF.pdfIMC_WFS_channels_TF.pdfIMC_WFS_segment_TF.pdfIMC_WFS_220804.xlsx
             Reply  Thu Aug 4 22:14:38 2022, Koji, Update, IOO, WFS investigation Screenshot_2022-08-04_22-17-19.pngScreenshot_2022-08-04_22-18-45.png
                Reply  Thu Aug 4 22:32:31 2022, Koji, Update, IOO, Upon leaving the lab (WFS investigation) 
                   Reply  Fri Aug 5 12:42:12 2022, Koji, Update, IOO, IMC WFS: Overnight observation  Screen_Shot_2022-08-05_at_12.04.01.png
Message ID: 17059     Entry time: Thu Aug 4 21:58:18 2022     In reply to: 17053     Reply to this: 17060
Author: Koji 
Type: Update 
Category: IOO 
Subject: WFS investigation 

OK... It seems that all the 6 dof of the IMC WFS servo loops were closed with some condition...

- Measured the transfer functions from ASCPIT/YAW_EXC of each suspensions to WFS segs.
- FInd the proper input matrix for PIT and YAW for WFS1 and WFS2
- Closed loops one by one => This was not so successful because the loop shape was quite conditional.
- Closed WFS1/WFS2 loops one by one only with FM4 (0.8Hz Zero / (100Hz pole)^2). Adjust the gains to have the UGF at a few Hz.

- Found that the separation between WFS1P and WFS2P was not good. This caused instability of these loops when the gains were matched. I ended up lowering the gain of WFS1P by a factor of 10. This made the loop OK to work. FM3 (Integrator below 0.8Hz) worked fine.

- FM9 Rolloff filters (RLP12) makes the loops unstable.

- The MC2 spot loops (MC2_TRANS_PIT/YAW) are supposed to be slow loops. From the time series behavior it looks they are working.


MEDM Snapshots (Attchaments 1~4)

Attachment 1: Screenshot_2022-08-04_22-10-57.png  48 kB  Uploaded Thu Aug 4 23:13:12 2022  | Hide | Hide all
Screenshot_2022-08-04_22-10-57.png
Attachment 2: Screenshot_2022-08-04_22-11-16.png  48 kB  Uploaded Thu Aug 4 23:13:21 2022  | Hide | Hide all
Screenshot_2022-08-04_22-11-16.png
Attachment 3: Screenshot_2022-08-04_22-11-53.png  48 kB  Uploaded Thu Aug 4 23:13:27 2022  | Hide | Hide all
Screenshot_2022-08-04_22-11-53.png
Attachment 4: Screenshot_2022-08-04_22-12-39.png  34 kB  Uploaded Thu Aug 4 23:13:34 2022  | Hide | Hide all
Screenshot_2022-08-04_22-12-39.png
ELOG V3.1.3-